# Berezin transform of Toeplitz operators on Bergman space with Békollé and Bonami weights

### v1

@@ 1. Introduction and Results
The pseudohyperbolic metric $d:\mathbb{D}\times\mathbb{D}\to[0,1)$ is defined as
@@
For $z\in\mathbb{D}$ and $r>0$, the pseudohyperbolic disk centered at $z$ with radius $r$ is given by
${\Delta(z,r)}:=\{w\in\mathbb{D}:d(z,w)<r\}.$
@@
where $u(\Delta(z,r))$ denotes the integral of the weight $u$ over the pseudohyperbolic disk $\Delta(z,r)$.
@@
Our main propose is to characterize some properties of Toeplitz operator $T_{\mu}:A_{u}^{q}\to A_{u}^{p}$ such as the boundedness, compactness and Schatten class in terms of its generalized t-Berezin transform and average function, and another condition related to the geometric characterization of the symbol $\mu$ to be Carleson measure for $A^{s}_{u}$ , for every possible cases of $s,p,q\in(0,\infty).$ The first two results give some equivalent necessary and sufficient conditions for the boundedness and compactness, respectively, of the operator $T_{\mu}$ acting between different Békollé and Bonami Bergman spaces $A_{u}^{q}(\mathbb{D})$ and $A_{u}^{q}(\mathbb{D}),$ for $0<p\leq q<\infty.$ Moreover, we get some estimates of its norm.
@@
The following Theorem gives s equivalent necessary and sufficient conditions for the boundedness(compactness) of the operator $T_{\mu},$ for $0<q<p<\infty.$
@@
where the constants involved depend only on $r$. Furthermore, we define a modified pseudohyperbolic disk which is compareable with the original one,
$\Delta^{\prime}(z,r):=\left\{w\in\mathbb{D}:\frac{|z-w|}{1-|w|}<r\right\},$
and $\Delta^{\prime}(z,r)\simeq\Delta(z,r)$.
We will also use the following class of weights which is denoted by condition $C_{p}$. For $p>1$, a positive locally integrable weights $u$ is said to belong to the $C_{p}$ class, (or satisfy condition $C_{p}$) if
@@
For $p>0$, the point evaluations on $A^{p}(u)$ are bounded linear functionals. To be precise, we have the following estimate.
@@
Within the Békollé-Bonami class of weights, Bergman metric balls with centers sufficiently close in the pseudohyperbolic metric have comparable weighted areas.
@@
Our method relies on a popular decomposition of $\mathbb{D}$ which is used repeatedly in many papers. See Theorem 1 in [15] for instance. And we will use the following covering lemma.
@@
In our main results, we used Carleson measures characterise the Toeplitz operator $T_{\mu}:A_{u}^{q}\to A_{u}^{p}$.
@@ 2.1. Carleson measures
Let $0<p,q<\infty$. A positive Borel measure $\mu$ on $\mathbb{D}$ is defined to be a $q$-Carleson measure for $A^{p}(u)$ if the embedding $I:A^{p}(u)\to L^{q}(\textrm{d}\mu)$ is bounded. We have the following Carleson embedding theorem.
@@
The equivalence (a) and (c) was proved by Constantin in [5]. And there also the compactness condition,
@@ 2.2. Kernel Estimates
Recall that $K(w,z)=K_{w}(z)$ denotes the reproducing kernels of $A_{u}^{2}$, and $k_{w}(z)=\frac{|K_{w}(z)|}{\|K_{w}\|_{A_{u}^{2}}}$ denotes the normalized reproducing kernel of $A_{u}^{2}$. In what follows we will introduce a $L^{p}$-estimate of $K_{w}(z)$ and also an estimate of $|k_{w}(z)|^{2}$.
@@
And recall thet the normalized reproducing kernel $k_{p,w}=\frac{K_{w}(z)}{\|K_{w}\|_{A_{u}^{p}}}$, by Lemma 2.1 and Lemma 2.6 above, we have
(2.7) $k_{p,w}\lesssim u(\Delta(w,r))^{-\frac{1}{p}}.$
And the following lemma estimates the reproducing kernel functions on the diagonal.
@@
Now we can estimate the normalized reproducing kernel $k_{w}(z)$ around the diagonal. Our approach adapts the methodology from Lemma 3.6 in [8] to fix the problems on the weighted Bergman spaces with Békollé weights.
@@
With Lemma 2.7 and Lemma 2.8, we can characterize the relationship between Berezin transform $\tilde{\mu}(z)$ and the average function $\widehat{\mu_{r}}(z)$.
@@
Conversely, by Lemma 2.1 one has that
@@
where $z\in\mathbb{D}$ and $r>0$. Let $\{a_{j}\}$ and $r>0$ be chosen as in Lemma 2.3, using Lemma 2.2, Fubini’s Theorem and Lemma 2.3, we can conduct the following computation,
@@
$=$ $\int_{\Delta(z,r)}\left|\frac{k_{z}\|K_{z}\|_{A_{u}^{2}}}{\|K(\cdot,z)\|_{A_{u}^{t}}}\right|^{t}d\mu(w)$
@@
$\lesssim$ ${\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z).$
@@ 2.3. Duality
(2.14) $\langle f,g\rangle_{A^{2}(u)}=\int_{\mathbb{D}}f(z)\overline{g(z)}u(z)\textrm{d}A(z).$
@@
For all $f\in A_{u}^{p^{\prime}}$ and $g\in A_{u}^{p}$, denote $\Lambda_{g}(f)=\langle f,g\rangle_{A_{u}^{2}}$. Then, using Hölder’s inequality, we have
$|\Lambda_{g}(f)|=|\langle f,g\rangle_{A_{u}^{2}}|\leq\int_{\mathbb{D}}|f(z)||g(z)|u(z)dA(z)\leq\|f\|_{A_{u}^{p^{\prime}}}\|g\|_{A_{u}^{p}},$
so $\Lambda_{g}\in(A_{u}^{p^{\prime}})^{*}$, and $\|\Lambda_{g}\|\leq\|g\|_{A_{u}^{p}}$
By the Hahn-Banach theorem, every bounded linear operator $T\in(A_{u}^{p^{\prime}})^{*}$ can be extended to a bounded linear functional $\widetilde{T}$ on $L_{u}^{p^{\prime}}$. Since $(A_{u}^{p^{\prime}})^{*}$ is isometrically isomorphic to $L_{u}^{p}$, there exists an $h\in L_{u}^{p}$ such that $T(f)=\langle f,h\rangle_{L_{u}^{2}}$ for all $f\in A_{u}^{p^{\prime}}$, and $\|\widetilde{T}\|=\|h\|_{L_{u}^{p}}$. Since the Bergman projection $P_{u}:L^{p}_{u}\to A^{p}_{u}$ is bounded and then using Fubini’s theorem, for any $f\in A_{u}^{p^{\prime}}$ we have
$T(f)=T(P_{u}f)=\langle f,P_{u}h\rangle_{A_{u}^{2}}=\Lambda_{g}(f),$
where $g=P_{u}h$. It follows that $g\in A_{u}^{p}$ and satisfies $\|g\|_{A_{i}^{p}}\leq\|P_{u}\|\|h\|_{L_{u}^{p}}=\|P_{u}\|\|T\|$. ∎
Using the duality above, we can prove the following lemma, which is essential for characterizing compactness properties.
@@
Let $p_{0}>1$, $0<p<\infty$. If $u\in B_{p_{0}}$, then the normalized kernel function $k_{p,w}(z)$ is weakly converge to zero as $|z|\to 1^{-}$.
@@
By Lemma 2.12, for $1<s<\infty$, it is suffices to show that
$|\langle k_{w},g\rangle_{A_{u}^{2}}|\to 0,\text{ as }|z|\to 1^{-},$
for all $g\in A_{u}^{s^{\prime}}$, where $\frac{1}{s}+\frac{1}{s^{\prime}}=1$. By the definition of $k_{w}$ and reproducing formula, we get
$|\langle k_{w},g\rangle_{A_{u}^{2}}|$ $=$ $\frac{\int_{\mathbb{D}}|K_{w}(z)||g(z)|u(z)dA(z)}{\|K_{w}\|_{A_{u}^{s}}}$
$=$ $\frac{|g(w)|}{\|K_{w}\|_{A_{u}^{s}}}.$
@@
$|\langle k_{w},g\rangle_{A_{u}^{2}}|\simeq\frac{|g(w)|(1-|z|)^{2}}{u(\Delta(w,r))^{\frac{1}{s}}},$
which converges to zero weakly as $|z|\to 1^{-}$. Then for any $0<p<\infty$, $k_{p,w}(z)=\frac{\|K_{z}\|_{A_{u}^{p}}}{\|K_{z}\|_{A_{u}^{2}}^{2}}k_{w}$, which also converges to zero weakly as $|z|\to 1^{-}$. ∎
The following Proposition is a classical result whose proof is similar to that given by K. Zhu in [17, Theorem 1.14].
@@
Let $p_{0}>1$ and $u\in B_{p_{0}}.$ A linear operator $T$ on $A^{2}(u)$ is compact if and only if $\|Tf_{n}\|_{A^{2}(u)}\longrightarrow 0$ whenever $f_{n}\to 0$ weakly in $A^{2}(u).$
3. case $0<p\leq q<\infty$
3.1. Proof of Theorem 1.1
@@
Suppose that Toeplitz operator $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is bounded. For case $q>1$, using Hölder’s inequality and the kernel estimate in Lemma 2.6, we have
@@
For case $q\leq 1$, by Lemma 2.8, inequality (2.7) and then Lemma 2.1, we have
$\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}$ $=$ $u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\int_{\mathbb{D}}|k_{z}(w)|^{2}d\mu(w)$
$\simeq$ $u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\int_{\mathbb{D}}\frac{|K(z,w)|^{2}}{\|K_{z}\|_{A_{u}^{2}}^{2}}d\mu(w)$
$\simeq$ $u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\frac{\|K_{z}\|_{A_{u}^{p}}}{\|K_{z}\|_{A_{u}^{2}}^{2}}\int_{\mathbb{D}}K(z,w)\frac{|K(z,w)|}{\|K_{z}\|_{A_{u}^{p}}}d\mu(w)$
$\simeq$ $u(\Delta(z,r))^{\frac{1}{q}}\int_{\mathbb{D}}K(z,w)k_{p,z}(w)\textrm{d}\mu(w)$
@@
$\lesssim$ $\|T_{\mu}k_{p,z}\|_{A_{u}^{q}}$
$\lesssim$ $\|T_{\mu}\|.$
In addition, by Corollary 2.11, we have $\frac{|\widetilde{\mu}_{t}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}$.
@@
$\int_{\mathbb{D}}|f(z)|d\mu(z)$ $\lesssim$ $\left(\int_{\mathbb{D}}|f(z)|^{{}^{\frac{pq^{\prime}}{p+q^{\prime}}}}u(z)dA(z)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}},$
which implies that $\mu$ is a 1-Carleson measure for $A_{u}^{\frac{pq^{\prime}}{p+q^{\prime}}}$, Using this fact together with Hölder’s inequality, for any analytic function $f\in A^{p}_{u}$ we obtain
@@
$\lesssim$ $\|f\|_{A_{u}^{p}}\|g\|_{A^{q^{\prime}}_{u}}.$
where $q\in A^{q^{\prime}}_{u}$ since duality (Lemma2.12). It follows that $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is bounded.
We now consider the case $q\leq 1$. Using some arguments like the case $q>1$, we can show that $\mu$ is also a 1-Carleson measure for $A_{u}^{\frac{pq^{\prime}}{p+q^{\prime}}}$. Since $q\leq 1$, the triangle inequality $(a+b)^{q}\leq a^{q}+b^{q}$ holds for non-negative $a$ and $b$. For $f\in A_{u}^{p}$ and $\{z_{k}\}_{k}$ is some $r$-lattice in Lemma 2.3, applying Lemma 2.1, we obtain

### v2

@@ 1. Introduction and Results
The pseudo-hyperbolic metric $d:\mathbb{D}\times\mathbb{D}\to[0,1)$ is defined as
@@
For $z\in\mathbb{D}$ and $r>0$, the pseudo-hyperbolic disk centered at $z$ with radius $r$ is given by
(2.1) ${\Delta(z,r)}:=\{w\in\mathbb{D}:d(z,w)<r\}.$
@@
where $u(\Delta(z,r))$ denotes the integral of the weight $u$ over the pseudo-hyperbolic disk $\Delta(z,r)$.
@@
Our main propose is to obtain characterizations of some properties of Toeplitz operator $T_{\mu}:A_{u}^{q}\to A_{u}^{p}$ such as the boundedness, compactness and Schatten classmembership in terms of its generalized t-Berezin transform and average function, and another condition related to the geometric characterization of the symbol $\mu$ to be Carleson measure for $A^{s}_{u},$ for every possible cases of $s,p,q\in(0,\infty).$ The first two results give some equivalent necessary and sufficient conditions for the boundedness and compactness, respectively, of the operator $T_{\mu}$ acting between different Békollé and Bonami Bergman spaces $A_{u}^{q}(\mathbb{D})$ and $A_{u}^{q}(\mathbb{D}),$ for $0<p\leq q<\infty.$
@@
The following theorem gives some equivalent necessary and sufficient conditions for the boundedness and compactness of Toeplitz operator $T_{\mu},$ for $0<q<p<\infty.$
@@
Moreover, in the last results of this paper, we characterize the essential norm of $T_{\mu}$ and its Schatten class membership.
Theorem 1.4.
$\|T_{\mu}\|_{e}\simeq\limsup_{|z|\to 1}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}\simeq\limsup_{|z|\to 1}\widehat{\mu_{r}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}.$
Theorem 1.5.
$\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)<\infty.$
@@
In this section we review a number of preliminary results related to the weight class, characterizations of Carleson measures and the Bergman kernels in the corresponding weighted spaces, which lead us to prove our main results.
@@
where the constants involved depend only on $r$. Furthermore, we define a modified pseudohyperbolic disk which is compareable with $\Delta(z,r)$(2.1),
$\Delta^{\prime}(z,r):=\left\{w\in\mathbb{D}:\frac{|z-w|}{1-|w|}<r\right\}.$
We will also use the following class of weights which is denoted by condition $C_{p}$. For $p>1$, a positive locally integrable weight $u$ belongs to the class $C_{p}$ (or satisfies condition $C_{p}$) if
@@
For $p>0$, the point evaluations on $A^{p}(u)$ are bounded linear functionals, as one can see from the following lemma.
@@
Within the Békollé-Bonami class of weights, Bergman metric balls with centers sufficiently close in the pseudo-hyperbolic metric have comparable weighted areas.
@@
Our method relies on a popular decomposition of $\mathbb{D}$ which is used repeatedly in many papers, e.g., Theorem 1 in [15]. We will use the following covering lemma.
@@ 2.1. Carleson measures
A positive Borel measure $\mu$ on $\mathbb{D}$ is a qq-(vanishing) Carleson measure for $A^{p}(u)$ if the embedding $I:A^{p}(u)\to L^{q}(\textrm{d}\mu)$ is (compact) bounded. The following two lemma give the characterizations of $q$- (vanishing) Carleson measure for $A^{p}(u),$ when $0<p\leq q$ and $u$ belongs to our class of weights.
To prove our main results we need some characterizations of the bounded and compact embedding operator $I,$ whenever $0<p,q<\infty,$ which are presented in the following lemmas.
@@
The equivalence (a) and (c) was proved by Constantin in [5].
@@ 2.2. Kernel Estimates
Recall that $K(w,z)=K_{w}(z)$ denotes the reproducing kernels of $A_{u}^{2}$, and $k_{w}(z)=\frac{|K_{w}(z)|}{\|K_{w}\|_{A_{u}^{2}}}$ denotes the normalized reproducing kernel of $A_{u}^{2}$. In what follows we will introduce a $L^{p}$-estimate of $K_{w}(z)$ and also a pointwise estimate of $|k_{w}(z)|^{2}$.
@@
And recall thet the normalized reproducing kernel $k_{p,w}=\frac{K_{w}}{\|K_{w}\|_{A_{u}^{p}}}$, by Lemma 2.1 and Lemma 2.6 above, we have
(2.7) $|k_{p,w}(z)|\lesssim u(\Delta(w,r))^{-\frac{1}{p}},\quad\textrm{for any}\,\,\,z,w\in\mathbb{D}.$
Next, the following lemma shows the estimates of the reproducing kernel functions on the diagonal.
@@
Now, the following lemma gives the pointwise estimates of $k_{w}(z),$ when $z$ and $w$ are closed.
@@
The remain Lemmas of this subsection show the relationship between Berezin transform $\tilde{\mu}_{t}(z)$ and the average function $\widehat{\mu_{r}}(z),$ for any $r,t>0.$
@@
Conversely, by Lemma 2.1, one has that
@@
where $z,w\in\mathbb{D}$ and $r>0$. Let $\{a_{j}\}$ and $r>0$ be chosen as in Lemma 2.3, using Lemma 2.2, Fubini’s Theorem and Lemma 2.3, we can conduct the following computation,
@@
$=$ $\int_{\Delta(z,r)}\left|\frac{k_{z}(w)\|K_{z}\|_{A_{u}^{2}}}{\|K_{z}\|_{A_{u}^{t}}}\right|^{t}d\mu(w)$
@@
$\lesssim$ ${\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z)\,\|k_{t,z}\|^{t}_{A^{t}_{u}}={\sup_{z\in\mathbb{D}}}\widehat{\mu_{r}}(z).$
@@
Furthermore,
$\tilde{\mu}(z)\simeq\tilde{\mu}_{t}(z)\simeq\widehat{\mu_{r}}(z).$
@@ 2.3. Duality
(2.14) $\langle f,g\rangle_{u}=\int_{\mathbb{D}}f(z)\overline{g(z)}u(z)\textrm{d}A(z).$
@@
For all $f\in A_{u}^{p^{\prime}}$ and $g\in A_{u}^{p}$, denote $\Lambda_{g}(f)=\langle f,g\rangle_{u}$. Then, using Hölder’s inequality, we have
$|\Lambda_{g}(f)|=|\langle f,g\rangle_{u}|\leq\int_{\mathbb{D}}|f(z)||g(z)|u(z)dA(z)\leq\|f\|_{A_{u}^{p^{\prime}}}\|g\|_{A_{u}^{p}},$
so $\Lambda_{g}\in(A_{u}^{p^{\prime}})^{*}$, and $\|\Lambda_{g}\|\leq\|g\|_{A_{u}^{p}}.$
By the Hahn-Banach theorem, every bounded linear operator $T\in(A_{u}^{p^{\prime}})^{*}$ can be extended to a bounded linear functional $\widetilde{T}$ on $L_{u}^{p^{\prime}}$. Since $(A_{u}^{p^{\prime}})^{*}$ is isometrically isomorphic to $L_{u}^{p}$, there exists an $h\in L_{u}^{p}$ such that $T(f)=\langle f,h\rangle_{u}$ for all $f\in A_{u}^{p^{\prime}}$, and $\|\widetilde{T}\|=\|h\|_{L_{u}^{p}}$. Since the Bergman projection $P_{u}:L^{p}_{u}\to A^{p}_{u}$ is bounded ([1, 2]) and then using Fubini’s theorem, for any $f\in A_{u}^{p^{\prime}}$ we have
$T(f)=T(P_{u}f)=\langle f,P_{u}h\rangle_{u}=\Lambda_{g}(f),$
where $g=P_{u}h$. It follows that $g\in A_{u}^{p}$ and satisfies $\|g\|_{A_{i}^{p}}\leq\|P_{u}\|\|h\|_{L_{u}^{p}}=\|P_{u}\|\|T\|$. This completes the proof. ∎
Then we can prove the following lemma, which is essential for characterizing compactness properties of operators acting on Békollé and Bonami Bergman spaces. Chacón [3] proved a version for $k_{2,w}$ case, we show it below for convenient.
Lemma 2.13 (Propositon 19 in [3]).
Let $p_{0}>1$. If $u\in B_{p_{0}}$, then the normalized kernel function $k_{p,w}$ is weakly converge to zero as $|w|\to 1^{-}$ in $A^{2}_{u}$.
Then we generalize Lemma 2.13 to arbitrary $p$-case.
@@
Let $p_{0}>1$, $1<p<\infty$. If $u\in B_{p_{0}}$, then the normalized kernel function $k_{p,w}$ is weakly converge to zero as $|w|\to 1^{-}$.
@@
By Lemma 2.12, for $1<p<\infty$, it suffices to show that
$|\langle k_{w},g\rangle_{u}|\to 0,\text{ as }|w|\to 1^{-},$
for all $g\in A_{u}^{p^{\prime}}$ and $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. By the definition of $k_{w}$ and reproducing kernel formula, we get
$|\langle k_{p,w},g\rangle_{u}|$ $=$ $\frac{1}{\|K_{w}\|_{A_{u}^{p}}}\Big|\int_{\mathbb{D}}K(z,w)\,g(z)\,u(z)\,dA(z)\Big|$
$=$ $\frac{|g(w)|}{\|K_{w}\|_{A_{u}^{p}}}.$
@@
$|\langle k_{p,w},g\rangle_{u}|\simeq\frac{|g(w)|(1-|w|)^{2}}{u(\Delta(w,r))^{\frac{1}{p}}},$
$|\langle k_{p,w},g\rangle_{u}|$ $\simeq$ $\frac{|g(w)|(1-|w|)^{2}}{u(\Delta(w,r))^{\frac{1}{p}}}$
$\lesssim$ $\frac{(1-|w|)^{2}}{u(\Delta(w,r))^{\frac{1}{p}}}\left(u(\Delta(w,r))^{-1}\int_{\Delta(w,r)}|g(z)|^{p^{\prime}}u(z)dA(z)\right)^{\frac{1}{p^{\prime}}}$
$\lesssim$ $\frac{(1-|w|)^{2}}{u(\Delta(w,r))}\|g\|_{A^{p^{\prime}}_{u}}$
$\simeq$ $\left|\langle k_{2,w},\frac{1}{u(\Delta(w,r))^{\frac{1}{2}}}\rangle_{u}\right|\|g\|_{A^{p^{\prime}}_{u}}$
then Lemma 2.13 implies that $k_{p,w}$ converges to zero weakly as $|w|\to 1^{-}.$
The following Proposition is a classical result and its proof is similar to one given by K. Zhu in [17, Theorem 1.14].
@@
Let $p_{0}>1$ and $u\in B_{p_{0}}.$ A linear operator $T$ on $A^{2}(u)$ is compact if and only if $\|Tf_{n}\|_{A^{2}(u)}\longrightarrow 0,$ whenever $f_{n}\to 0$ weakly in $A^{2}(u).$
3. Proof of Theorem 1.1
@@
For the case $q\leq 1$, by Lemma 2.8, inequality (2.7) and then Lemma 2.1, we have
$\Big|\frac{\widetilde{\mu}(z)}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\Big|$ $=$ $\Big|u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\int_{\mathbb{D}}|k_{z}(w)|^{2}d\mu(w)\Big|$
$\simeq$ $\Big|u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\int_{\mathbb{D}}\frac{|K(z,w)|^{2}}{\|K_{z}\|_{A_{u}^{2}}^{2}}d\mu(w)\Big|$
$\simeq$ $\Big|u(\Delta(z,r))^{\frac{1}{q}-\frac{1}{p}}\frac{\|K_{z}\|_{A_{u}^{p}}}{\|K_{z}\|_{A_{u}^{2}}^{2}}\int_{\mathbb{D}}K(z,w)\frac{\overline{K(z,w)}}{\|K_{z}\|_{A_{u}^{p}}}d\mu(w)\Big|$
$\simeq$ $\Big|u(\Delta(z,r))^{\frac{1}{q}}\int_{\mathbb{D}}K(z,w)k_{p,z}(w)\textrm{d}\mu(w)\Big|$
@@
$\lesssim$ $\|T_{\mu}k_{p,z}\|_{A_{u}^{q}}\leq\|T_{\mu}\|.$
Thus,
$\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}.$
In addition, by Corollary 2.11, we have $\frac{|\widetilde{\mu}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\simeq\frac{|\widetilde{\mu}_{t}(z)|}{u(\Delta(z,r))^{\frac{1}{p}-\frac{1}{q}}}\in L^{\infty}$, for any $t>0.$
@@
$\int_{\mathbb{D}}|h(z)|\,d\mu(z)$ $\lesssim\left(\int_{\mathbb{D}}|h(z)|^{{}^{\frac{pq^{\prime}}{p+q^{\prime}}}}u(z)dA(z)\right)^{\frac{p+q^{\prime}}{pq^{\prime}}},$
for any analytic function $f$ which could have zero points on $\mathbb{D}$, we consider $f_{\varepsilon}=f+\varepsilon,\varepsilon>0$, which is also analytic without zeros on $\mathbb{D}.$ By Fatou’s Lemma, and (3) and dominated convergence theorem, we get
$\int_{\mathbb{D}}|f|d\mu=\lim_{\varepsilon\to 0}\int_{\mathbb{D}}|f_{\varepsilon}|d\mu\leq\liminf_{\varepsilon\to 0}C\left(\int_{\mathbb{D}}|f_{\varepsilon}|^{\frac{pq^{\prime}}{p+q^{\prime}}}u\,dA\right)^{\frac{p+q^{\prime}}{pq^{\prime}}}\lesssim\left(\int_{\mathbb{D}}|f|^{\frac{pq^{\prime}}{p+q^{\prime}}}u\,dA\right)^{\frac{p+q^{\prime}}{pq^{\prime}}},$
which implies that $\mu$ is a 1-Carleson measure for $A_{u}^{\frac{pq^{\prime}}{p+q^{\prime}}}.$ This together with applying Hölder’s inequality, for any analytic function $f\in A^{p}_{u},$ we obtain
@@
$\lesssim$ $\|f\|_{A_{u}^{p}}\|g\|_{A^{q^{\prime}}_{u}},$
where $g\in A^{q^{\prime}}_{u}$ since duality (Lemma 2.12). It follows that $T_{\mu}:A^{p}_{u}\to A^{q}_{u}$ is bounded.
We now consider the case $q\leq 1$. Using some arguments like in the case $q>1$, we can show that $\mu$ is also a 1-Carleson measure for $A_{u}^{\frac{pq^{\prime}}{p+q^{\prime}}}$. Since $q\leq 1$, the triangle inequality $(a+b)^{q}\leq a^{q}+b^{q}$ holds for non-negative $a$ and $b$. For $f\in A_{u}^{p}$ and $\{z_{k}\}_{k}$ is a $r$-lattice as defined in Lemma 2.3. By Lemma 2.1, we obtain